OpenStudy (anonymous):

solve (2.8 + 0.1 * x / 8 + x ) = 25/100 @mathstudent55

OpenStudy (anonymous):

@immy79

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

\[\left(\frac{28}{10}+\frac{1}{10}*\frac{x}{8}+x\right)=\frac{25}{100} \]Is this equivalent to the problem expression?

OpenStudy (anonymous):

no its 2.8 not 28

OpenStudy (anonymous):

Well if you dividr 28 by 10 you get 2.8 . Why eliminate decimal numbers? On average the solution will calculate to an exact symbolic answer.

OpenStudy (anonymous):

oh okay so the problem looks okay now

OpenStudy (anonymous):

Mathematica says that \[x=-\frac{68}{27} \]The above computed to 20 decimal digits. -2.5185185185185185185

OpenStudy (anonymous):

Sorry about the word dividr. That is the word divide, misspelled.

OpenStudy (anonymous):

I''m confused how did you get that?

OpenStudy (anonymous):

Working on it.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\left(\frac{28}{10}+\frac{1}{10}*\frac{x}{8}+x\right)\text{=}\frac{25}{100}\]Move the 28/10 to the RHS.\[\left(\frac{1}{10}*\frac{x}{8}+x\right)\text{=}\frac{25}{100}-\frac{28}{10}\]Collect the terms on both sides and simplify.\[\left(\frac{81 x}{80}\right)\text{= }-\frac{51}{20}\]Multiply both equation sides by 80/81 to get ride of the coefficient of x and then simplify the result on the RHS.\[\left(\frac{80}{81}\right)\left(\frac{81 x}{80}\right)\text{=}-\frac{51}{20}\left(\frac{80}{81}\right)\]\[x\text{=}-\frac{68}{27}\]

OpenStudy (anonymous):

oh ok thank you.

OpenStudy (anonymous):

Your welcome and thank you for the medal.